Find residue
1)- 4/1+z^2
2)- 1/1-e^z
3)- sin2z/z^6
1
Expert's answer
2020-10-12T13:32:10-0400

"\displaystyle(1) \\\textsf{Residue of} \hspace{0.1cm} f_1(z) = \frac{1}{z^2 + 1} \\

\displaystyle f(z)\hspace{0.1cm}\textsf{has two simple poles at}\hspace{0.1cm}z = \pm j ,\textsf{where}\hspace{0.1cm} j \hspace{0.1cm}\textsf{is a complex number}\\
\textsf{Calculating the residues at each poles.}\\
\textsf{The residue at} \hspace{0.1cm}z = j \hspace{0.1cm}\textsf{is}\\
\begin{aligned}
\lim_{z \rightarrow j} \left(\frac{1}{(z – j)(z + j)}\cdot (z – j)\right) &= \lim_{z \rightarrow j} \left(\frac{1}{z+j}\right) \\
&= \lim_{z \rightarrow j} \frac{1}{2j} = \frac{1}{2j} = \frac{-j}{2}
\end{aligned}\\

\displaystyle\textsf{The residue at} \hspace{0.1cm} z = -j \hspace{0.1cm}\textsf{is}\\
\begin{aligned}
\lim_{z \rightarrow -j} \left(\frac{1}{(z – j)(z + j)}\cdot (z – j)\right) &= \lim_{z \rightarrow -j} \left(\frac{1}{z - j}\right) 
\\&= \lim_{z \rightarrow -j} \frac{1}{-2j} = \frac{1}{-2j} = \frac{j}{2}
\end{aligned} \\



(2)\\
\displaystyle\textsf{Residue of} \hspace{0.1cm} f_2(z) = \frac{1}{1 - e^z} \\
f_2(z)\hspace{0.1cm}\textsf{has a pole when}\hspace{0.1cm} e^z = 1,\\ 
\implies z = 2jn\pi, \hspace{0.2cm} \forall n \in \mathbb{Z} \\
\therefore f_2(z)\hspace{0.1cm}\textsf{has poles at}\hspace{0.1cm} z = 2jn\pi\\
\textsf{Deriving the Laurent's series of}\hspace{0.1cm} f_2(z) \\
\displaystyle\begin{aligned}
e^z - 1 &=\sum_{n=0}^\infty \frac{z^n}{n!} - 1 = \sum_{n=1}^\infty \frac{z^n}{n!} \\
&= z\sum_{n=1}^\infty \frac{z^{n - 1}}{n!} = z\sum_{n=0}^\infty \frac{z^n}{(n + 1)!} \\
&= z\left(1 + \sum_{n=1}^\infty \frac{z^n}{(n + 1)!}\right)
\end{aligned} \\
\begin{aligned}
\frac{1}{e^z - 1} &= \frac{1}{z}\left(1 + \sum_{n=1}^\infty \frac{z^n}{(n + 1)!}\right)^{-1}
\\&= \frac{1}{z}\sum_{k=0}^\infty (-1)^k \left(\sum_{n=1}^\infty \frac{z^n}{(n + 1)!}\right)^k
\\&= \frac{1}{z} - \frac{1}{z}\left(\frac{z}{2!} + \frac{z^2}{3!} +...\right)
\\&+ \frac{1}{z}\left(\frac{z}{2!} + \frac{z^2}{3!} +...\right)^2 \\&- \frac{1}{z}\left(\frac{z}{2!} + \frac{z^2}{3!} +...\right)^3 
\\&+ \frac{1}{z}\left(\frac{z}{2!} + \frac{z^2}{3!} +...\right)^4+...
\\&= \frac{1}{z} - \frac{1}{2!} + \left(-\frac{1}{3!} + \frac{1}{(2!)^2}\right)z 
\\&+ \left(-\frac{1}{4!} + \frac{2}{2! \cdot 3!} - \frac{1}{(2!)^3}\right)z^2 
\\&+ \left(-\frac{1}{5!} + \frac{2}{2! \cdot 4!} + \frac{1}{(3!)^2} - \frac{3}{(2!)^2\cdot 3!} + \frac{1}{(2!)^4}\right) z^3 +...
\\&= \frac{1}{z} - \frac{1}{2} + \frac{z}{12} - \frac{z^3}{720} +...
\end{aligned}\\

\displaystyle\textsf{By Cauchy's theorem,}\hspace{0.1cm} \frac{1}{2j\pi} \int_C \frac{1}{z^n} \, \mathrm{d}z =
\begin{cases} 
=0, &\textsf{for}\hspace{0.1cm} n = 2, 3, 4,...\\
= 2\pi j, &\textsf{for}\hspace{0.1cm} n= 1\\
= 0,&\textsf{for}\hspace{0.1cm} n = -1,-2,-3,-4,...
\end{cases} \\

\displaystyle\begin{aligned}
\therefore Res\left(\frac{1}{1 - e^z}, z = 2jn\pi\right) 
\\&= -\frac{1}{2j\pi}\int_C \left(\frac{1}{z} - \frac{1}{2} + \frac{z}{12} - \frac{z^3}{720} +...\right) \mathrm{d}z\\&= -\frac{2 \pi j}{2 \pi j} = - 1
\end{aligned} \\

\displaystyle(3) \\\textsf{Residue of} \hspace{0.1cm} f_3(z) = \frac{\sin(2z)}{z^6} \\
f(z)\hspace{0.1cm}\textsf{has a pole at}\hspace{0.1cm} z = 0\\
\textsf{The Taylor series of}\hspace{0.1cm} \sin(2z) \hspace{0.1cm} \textsf{is}\\
\sin{(2z)} = (2z) - \frac{(2z)^3}{3!} + \frac{(2z)^5}{5!} -\frac{(2z)^7}{7!}+ O(z^9)\\
\frac{\sin(2z)}{z^6}= \frac{2}{z^5} - \frac{8}{6z^3} + \frac{4}{15z} - \frac{8z}{315} + O(z^3)\\

Res\left(\frac{1}{z^n}, z=0\right) = \frac{1}{2j\pi} \int_C \frac{1}{z^n} \, \mathrm{d}z \hspace{0.1cm} \textsf{By definition} \\

\textsf{By Cauchy's theorem,}\hspace{0.1cm} \frac{1}{2j\pi} \int_C \frac{1}{z^n} \, \mathrm{d}z =

\begin{cases} 
=0, &\textsf{for}\hspace{0.1cm} n = 2, 3, 4,...\\
= 2\pi j, &\textsf{for}\hspace{0.1cm} n= 1\\
= 0,&\textsf{for}\hspace{0.1cm} n = -1,-2,-3,-4,...
\end{cases}\\
\therefore Res(f_3(z), z=0)) = 0 - 0 + \frac{4}{15}\cdot\frac{2\pi j}{2 \pi j} - 0 + 0 - 0 = \frac{4}{15}"